# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v3

@@ 7.2. LLMs validate response quality less reliably than crowdworkers
Validation in crowdsourcing workflows often involves asking a different worker to verify if errors or desirable qualities exist in outputs. However, both the literature and our case studies show mixed results on the effectiveness of this strategy for chaining.
While some studies show that prompting a model to reflect can help achieve higher accuracy outputs (Madaan et al., 2023b; Xie et al., 2023; Gero et al., 2023; Yao et al., 2023), others caution against the same model validating its own outputs, as LLMs exhibit high confidence for incorrect prior outputs (Liang et al., 2023; Du et al., 2023). Some approaches for more robust validation include: validating with a different LLM (Liang et al., 2023), debate-based communicative structures (Liang et al., 2023; Du et al., 2023), and structural validation such as similarity across parallel redundant streams (Liem et al., 2011). One paper found that “ease of persuasion,” or how quickly an LLM relented in a debate, was more indicative of factual accuracy than other measures of confidence (Du et al., 2023).
@@
Validation should address specific elements of quality that are aligned to worker capabilities. In our case studies, we determined what elements of quality to validate by evaluating intermediate outputs in a trial and error process. Determining these qualities can be tricky, especially as the capabilities of LLMs rapidly change. To determine what elements of quality require validation, intuition needs to be tempered by empirical testing and subtask evaluation.
Even after the set of qualities under consideration has been determined, we found LLMs could still struggle with successfully validating these qualities. To adapt to these issues, we used several approaches. First, when possible we made evaluate steps programmatic rather than LLM-based. This approach is most reliable, but there are a limited number of domains where programmatic checks are feasible. Second, we continued to break qualities down into even simpler parts when needed to align subtask difficulty with worker capabilities. Third, we simplify context to the minimum needed. Considering the story suggestion quality checks above, we found that including multiple prior suggestions, or including the entire story when not necessary, reduced the quality of validation checks.
@@
Some tasks have a large space of possible correct outputs, due to uncertainty or creativity, and could benefit from the user giving direct input. Under these conditions, it is helpful to draw on longstanding work on mixed-initiative user interfaces (Horvitz, 1999) to incorporate users into the workflow. Such an approach allows for parts of the workflow to be automated while others involve user input.
@@
User control through prompting is brittle (Arora et al., 2022), so workflows can enable more fine grained control. LLM chains can draw from user interaction methods from crowdsourcing workflows. Several large types of user interaction emerge from the literature.
@@
Third, a user may review options between iterations of the workflow (Willett et al., 2012; Cheng and Bernstein, 2015; Kulkarni et al., 2014; Park et al., 2022; Mirowski et al., 2023). This approach can help with exploring design spaces (Park et al., 2022; Kulkarni et al., 2014; Liu et al., 2023) and determining when the final quality threshold is met (Willett et al., 2012).
Fourth, a user may directly manipulate the output to see the breadth of user responses and to make final decisions about the output with less effort than doing the task directly (Chang et al., 2017; Chilton et al., 2013; Willett et al., 2012; Cheng and Bernstein, 2015; Lee et al., 2017; Mohanty et al., 2019; Bernstein et al., 2010). This approach enables users to surface uncertainty in the data and express control over the final output.
Many chaining papers discuss a generic possibility of a user replacing steps or intervening if needed (Schick et al., 2022; Yang et al., 2022; Wu et al., 2023a; Wang et al., 2022; Wu et al., 2023b; Zhang et al., 2024), but few actually implement this with people (Park et al., 2022; Kim et al., 2023b; Mirowski et al., 2023). We can see from crowdsourcing that there are many ways of incorporating the user that bring myriad benefits. In our case studies, we take direct manipulation from crowdsourcing and apply it to chaining, increasing the controllability of LLM outputs.
@@ 7.4. On-demand availability of LLMs reduces barriers to real-time crowdsourcing architectures
Methods like retainer models that incentivize keeping crowds were important for crowdsourcing workflows that involve real-time interaction for users (Bernstein et al., 2011) and for communicative workflows like debate in which workers must remain online as other workers respond to them (Drapeau et al., 2016; Chen et al., 2019). These workflows improve latency for users and improve quality and accuracy, respectively, but the challenges of retaining workers increased barriers in crowdsourcing. Similarly, for mixed-initiative interfaces, delayed worker responses increased the latency of interactions (Lasecki et al., 2015). Reduced latency of worker responses can make user interactions more real-time.
In other workflows, such as Flash Organizations (Valentine et al., 2017), workers recruit other workers to complete subtasks that are dynamically created. Finding the right workers for the job took a median of 14 minutes with a light screening process and of 15 hours with a more intense screening process (Valentine et al., 2017). In contrast, the Orchestrator LLM of the AutoGen (Wu et al., 2023a) chain “recruits” with other calls to LLMs, which can happen flexibly and quickly (Wu et al., 2023a).
@@
7.5. LLMs need different, often simpler, subtasks
Throughout our case studies, we often needed to break crowdsourcing subtasks into multiple, more specified subtasks. Multiple factors necessitated these changes: natural language is difficult to structure (e.g., thus, we query for one item category at a time rather than eight in Cascade), LLMs have difficulty balancing multiple requirements in a subtask (Wu et al., 2023c; Madaan et al., 2023b; Hsieh et al., 2023) (e.g., in response, we split Soylent’s “Fix” step into separate edit, merge, and delete queries), and LLMs do not understand length-related constraints (e.g., to address this, we add a hard-coded validation step after a generate subtask in Mechanical Novel to check for length).
In our case studies, LLMs performed better on highly specified subtasks for chains rather than fewer, more complicated subtasks. However, this may change as newer, more powerful LLMs are introduced.
@@ 7.6. Instructions should provide the minimum necessary context
Providing global context to subtasks is widely agreed to be a critical element of subtask design and output coherence, but it can increase instruction complexity (Wu et al., 2022a; Salehi et al., 2017; Ambati et al., 2012; Luther et al., 2015). However, findings from the literature survey and case studies suggest that LLMs can respond poorly to long, complicated instructions (Du et al., 2023; Parameswaran et al., 2023; Madaan et al., 2023b; Wu et al., 2023c). This context limitation should be re-evaluated over time as LLM context windows grow.
@@ 8. Future opportunities for chaining
8.1. Creative tasks with dynamic and communicative architectures
Crowdsourcing studies currently support a wider range of tasks for creative work with a wider range of approaches than chaining studies. We expect these approaches to transfer to chaining and bring myriad benefits.
Future work could attempt creative tasks that are under-explored in chaining, such as: prototyping (Lee et al., 2017; Valentine et al., 2017; Kulkarni et al., 2014), categorizing (Huffaker et al., 2020; Bragg et al., 2013; Chang et al., 2017), topic concept maps (Liu et al., 2018; Hahn et al., 2016; Luther et al., 2015), project development (Mahyar et al., 2018), or more specific tasks like itinerary creation (Zhang et al., 2012), researching a purchase (Kittur et al., 2011), and conceptual blends (Chilton et al., 2019).
Crowdsourcing work also uses a broader range of communicative and dynamic architectures to support creative work. Giving workers roles during communication can assist both crowdsourcing workflows (Valentine et al., 2017; Agapie et al., 2015; Retelny et al., 2014) and LLM chains (Wu et al., 2023a). In crowdsourcing work, some studies go beyond giving a role in instructions and instead create free-form role-based (Nebeling et al., 2016; Valentine et al., 2017; Zhang et al., 2023b), hierarchical (Agapie et al., 2015; Retelny et al., 2014), and state-sharing (Lasecki et al., 2015; Lee et al., 2017; Nebeling et al., 2016; Kulkarni et al., 2014; Mahyar et al., 2018; Lasecki et al., 2013) communicative architectures.
Dynamic architectures provide more flexibility when contingencies arise (Kim and Monroy-Hernandez, 2016; Valentine et al., 2017; Kittur et al., 2011; Kulkarni et al., 2012) and adapt to edge cases (Kim et al., 2014; Huang et al., 2015; Park et al., 2022; Kim et al., 2023b; Schick et al., 2022; Creswell and Shanahan, 2022; Yao et al., 2023; Zhou et al., 2022b; Nair et al., 2023b; Kulkarni et al., 2012; Hahn et al., 2016; Kittur et al., 2012; Wu et al., 2023c, 2022a). LLM chains rarely implement dynamic architectures for creative work (Yao et al., 2023; Wu et al., 2023c; Kim et al., 2023b; Zhang et al., 2023a). Crowdsourcing work applies dynamic architectures to a wider variety of creative tasks (Valentine et al., 2017; Lasecki et al., 2015; Retelny et al., 2014; Chilton et al., 2019; Kittur et al., 2011) and includes more complex dynamic workflows such as multi-level map-reduce (Kittur et al., 2011) and dynamically choosing between workflows using their maximum expected value of utility (Goto et al., 2016). Decision theoretic techniques, such as Partially Observable Markov Decision Processes, yield great efficiencies in crowdsourcing (Dai et al., 2010, 2013), but have thus far barely been considered for LLM chaining (Madaan et al., 2023a).
We recommend pursuing a broader range of creative tasks with more communicative (e.g. establishing a hierarchy of roles) and dynamic (e.g. multi-level map reduce) architectures.
@@ 8.2. Integrating the user
Although many chaining papers discuss involving humans into the chain hypothetically, few implement this user oversight (Stretcu et al., 2023; Toubal et al., 2024). Our case studies showcase one example of interactions through directly manipulable outputs, but the varied ways of incorporating the user from crowdsourcing workflows could continue to inspire similar approaches in chaining.
We recommend incorporating more approaches to user guidance from crowdsourcing, such as explorable interfaces (Willett et al., 2012; Chang et al., 2017; Cheng and Bernstein, 2015; Mohanty et al., 2019), worker-initiated revision (Salehi et al., 2017), and guided prototyping (Nebeling et al., 2016; Lasecki et al., 2015). We also recommend research into the the opportunities and limitations of user guidance, especially for creative tasks, as initial findings conflict on if using generative AI tools leads to divergent thinking (chi, 2023) or design fixation (Wadinambiarachchi et al., 2024).
@@ 8.3. Studying chains more systematically
More systematic studies on optimal chain design would be valuable. Several studies already approach questions around model choice (Madaan et al., 2023a; Du et al., 2023), incorporating roles into prompts (Wu et al., 2023a; Park et al., 2022), and using self-reflection or debate to validate outputs (Du et al., 2023; Liang et al., 2023). Although the outcomes may be different, systematic crowdsourcing studies could provide inspiration for some of these explorations. Prior crowdsourcing studies investigate the effect of context timing, architecture, and content on task output (Luther et al., 2015; Salehi et al., 2017). Others measure tradeoffs of iterative and parallel workflows (e.g., effect on latency, quality averages and variance, risk of cascading errors) in various task contexts, as well as good practices for each (e.g., appropriate number of branches or iterations and what context to share amongst branches) (Kittur et al., 2011; Chang et al., 2017; Luther et al., 2015; Huffaker et al., 2020; Little et al., 2010a; Goto et al., 2016; André et al., 2014).
By specifying a strategy of interest, the design space can orient studies for understanding optimal tactical design decisions. For example, studying debate designs could look at debate to establish adaptability for tasks with many edge cases, debate to validate outputs, or debate to encourage a diversity of creative ideas. Evaluating the use of tactic designs within the context of specific strategies will lead to more generalizable findings than evaluating solely on task performance.

### v4

@@ 8.2. LLMs validate response quality less reliably than crowdworkers
Validation in crowdsourcing workflows often involves asking a different actor to verify if errors or desirable qualities exist in outputs. However, both the literature and our case studies show mixed results on the effectiveness of this strategy for chaining.
While some studies show that prompting a model to reflect can help achieve higher accuracy outputs for objective tasks (Madaan et al., 2024; Xie et al., 2023; Gero et al., 2023; Yao et al., 2024), others caution against the same model validating its own outputs, as LLMs exhibit high confidence for incorrect prior outputs (Liang et al., 2023; Du et al., 2023). Some approaches for more robust validation include: validating with a different LLM (Liang et al., 2023), debate-based communicative structures (Liang et al., 2023; Du et al., 2023), and structural validation such as similarity across parallel redundant streams (Liem et al., 2011). One paper found that the “ease of persuasion,” or how quickly an LLM relented in a debate, was more indicative of factual accuracy than other measures of confidence (Du et al., 2023).
@@
Validation should address specific elements of quality that are aligned with actor capabilities. In our case studies, we determined what elements of quality to validate by evaluating intermediate outputs in a trial and error process. Determining these qualities can be tricky, especially as the capabilities of LLMs rapidly change. To determine what elements of quality require validation, intuition needs to be tempered by empirical testing and subtask evaluation.
Even after the set of qualities under consideration has been determined, we found LLMs could still struggle with successfully validating these qualities. To adapt to these issues, we used several approaches. First, when possible we made evaluate steps programmatic rather than LLM-based. This approach is most reliable, but there are a limited number of domains where programmatic checks are feasible. Second, we continued to break qualities down into even simpler parts when needed to align subtask difficulty with actor capabilities. Third, we simplify context to the minimum needed. Considering the story suggestion quality checks above, we found that including multiple prior suggestions, or including the entire story when not necessary, reduced the quality of validation checks.
@@
Some tasks have a large subjective space of possible correct outputs, due to uncertainty or creativity, and could benefit from the user giving direct input. Under these conditions, it is helpful to draw on longstanding work on mixed-initiative user interfaces (Horvitz, 1999). Such an approach allows for parts of the workflow to be automated while others involve user input.
@@
User control through prompting is brittle (Arora et al., 2023), so workflows can enable more fine grained control. LLM chains can draw from user interaction methods from crowdsourcing workflows. Several large types of user interaction emerge from the literature.
@@
Third, a user may review options between iterations of the workflow (Cheng and Bernstein, 2015; Mirowski et al., 2023). This approach can help with exploring design spaces (Park et al., 2022; Kulkarni et al., 2014; Liu et al., 2023) and determining when the final quality threshold is met (Willett et al., 2012).
Fourth, a user may directly manipulate the output to see the breadth of user responses and to make final decisions about the output with less effort than doing the task directly (Chang et al., 2017; Mohanty et al., 2019; Bernstein et al., 2010; Cheng and Bernstein, 2015). This approach enables users to surface uncertainty in the data and express control over the final output.
Many chaining papers discuss a generic possibility of a user replacing steps or intervening if needed (Schick et al., 2022; Yang et al., 2022; Wang et al., 2022; Wu et al., 2024), but few actually implement this with people (Park et al., 2022; Kim et al., 2023b; Mirowski et al., 2023). We can see from crowdsourcing that there are many ways of incorporating the user that bring myriad benefits. In our case studies, we take direct manipulation from crowdsourcing and apply it to chaining, increasing the controllability of LLM outputs.
@@ 8.4. On-demand availability of LLMs reduces barriers to real-time crowdsourcing architectures
Methods like retainer models that incentivize keeping crowds were important for crowdsourcing workflows that involve real-time interaction for users (Bernstein et al., 2011) and for communicative workflows like debate in which actors must remain online as other actors respond to them (Drapeau et al., 2016; Chen et al., 2019). These workflows improve latency for users and improve quality and accuracy for objective tasks, respectively, but the challenges of retaining crowdworkers increased barriers in crowdsourcing. Similarly, for mixed-initiative interfaces, delayed crowdworker responses increased the latency of interactions (Lasecki et al., 2015). Reduced latency of worker responses can make user interactions more real-time.
In other workflows, such as Flash Organizations (Valentine et al., 2017), crowdworkers recruit other crowdworkers to complete subtasks that are dynamically created. Finding the right crowdworkers for the job took a median of 14 minutes with a light screening process and 15 hours with a more intense screening process (Valentine et al., 2017). In contrast, the Orchestrator LLM of the AutoGen (Wu et al., 2023a) chain “recruits” with other calls to LLMs, which can happen flexibly and quickly (Wu et al., 2023a).
@@ 8.5. Instructions should provide the minimum necessary context
Providing global context to subtasks is widely agreed to be a critical element of subtask design and output coherence, but it can increase instruction complexity (Wu et al., 2022a; Salehi et al., 2017; Ambati et al., 2012; Luther et al., 2015). Crowdsourcing research more broadly has studied the process of improving worker instruction, including by ensuring comprehension through tests (Liu et al., 2016) and communicating misconceptions to requesters (Bragg et al., 2018). However, findings from the literature survey and case studies suggest that LLMs can respond poorly to long, complicated instructions (Du et al., 2023; Parameswaran et al., 2023; Madaan et al., 2024; Wu et al., 2023b). This context limitation should be re-evaluated over time as LLM context windows grow.
@@
8.6. LLMs need different, often simpler, subtasks
Throughout our case studies, we often needed to break crowdsourcing subtasks into multiple, more specified subtasks. Multiple factors necessitated these changes: natural language is difficult to structure (e.g., thus, we query for one item category at a time rather than eight in Cascade), LLMs have difficulty balancing multiple requirements in a subtask (Wu et al., 2023b; Madaan et al., 2024; Hsieh et al., 2023) (e.g., in response, we split Soylent’s “Fix” step into separate edit, merge, and delete queries), and LLMs do not understand length-related constraints (e.g., to address this, we add a hard-coded validation step after a generate subtask in Mechanical Novel to check for length).
In our case studies, LLMs performed better on highly specified subtasks for chains rather than on fewer, more complicated subtasks. However, this may change as newer, more powerful LLMs are introduced.
8.7. Chains can be useful despite increased cost
Although chains bring many improvements over zero-shot performance in our case studies, they also incur sizeable increases in expense and latency. Without thoughtful design, AI support tools can make easy tasks easier and hard tasks harder (Simkute et al., 2024). When should one choose between chaining or a zero-shot approach for potential applications? We consider the effects of cognitive load, the degree of user agency, the interruptibility of the overall workflow, and limitations in model capabilities.
The increased effort and latency of a chain’s runtime can be offset by quicker and less effortful interpretation of the outputs. For instance, in writing editing, it is easier to review highlighted minor edits rather than entirely new large blocks of text.
Chains may be helpful, despite higher costs, when the goal is spurring users’ innovation, learning, and design, rather than giving responses outright. For example, Metaphorian allows users to create scientific metaphors in an interactive interface (Kim et al., 2023b). The goal of such a system is not only to output the metaphors but also to be exploratory and spur thinking. Compared to zero-shot creation of metaphors, the paper finds users retain a greater sense of agency, metaphors increase in originality and coherence, and user satisfaction is higher when using an interactive chain.
The impact of increased latency on the utility of chains lessens when the broader workflow can be interrupted without major disruption. Chains that do not require user guidance, or only require it in spaced-out intervals, can run in the background while the user performs other tasks. Some tasks, such as in-line coding assistance, are not conducive to this type of interruption. However, others, such as short-story generation, can incorporate a long runtime, thus reducing the impact of longer latency on the user experience.
Finally, another reason to use chains despite the cost is if the task cannot be completed without a chain. At the moment, long-form writing and generating large code repositories are outside the capabilities of models and require a chain. It is worth noting that these capabilities will change over time. Thus, thoughtfulness about objectives and the broader task context can motivate the use of chains when a higher utility offsets increased costs.
@@ 9. Future opportunities for chaining
9.1. Subjective tasks with dynamic and communicative architectures
Crowdsourcing studies currently support a wider range of tasks for subjective work with a wider range of approaches than chaining studies. We expect these approaches to transfer to chaining and bring myriad benefits.
Future work could attempt subjective tasks that are under-explored in chaining, such as: prototyping (Lee et al., 2017; Valentine et al., 2017; Kulkarni et al., 2014), categorizing (Huffaker et al., 2020; Bragg et al., 2013; Chang et al., 2017), topic concept maps (Liu et al., 2018; Hahn et al., 2016; Luther et al., 2015), project development (Mahyar et al., 2018), or more specific tasks like itinerary creation (Zhang et al., 2012), researching a purchase (Kittur et al., 2011), and conceptual blends (Chilton et al., 2019).
Crowdsourcing work also uses a broader range of communicative and dynamic architectures to support subjective work. Giving actors roles during communication can assist both crowdsourcing workflows (Valentine et al., 2017; Agapie et al., 2015; Retelny et al., 2014) and LLM chains (Wu et al., 2023a; Li et al., 2024). In crowdsourcing work, some studies go beyond giving a role in instructions and instead create free-form role-based (Nebeling et al., 2016; Valentine et al., 2017; Zhang et al., 2023b), hierarchical (Agapie et al., 2015; Retelny et al., 2014), and state-sharing (Lasecki et al., 2015; Lee et al., 2017; Nebeling et al., 2016; Lasecki et al., 2013) communicative architectures.
Dynamic architectures provide more flexibility when contingencies arise (Kim and Monroy-Hernandez, 2016; Valentine et al., 2017; Kittur et al., 2011; Kulkarni et al., 2012) and adapt to edge cases (Nair et al., 2023b; Kulkarni et al., 2012; Chen et al., 2019). LLM chains rarely implement dynamic architectures for subjective creative work (Yao et al., 2024; Wu et al., 2023b; Kim et al., 2023b; Zhang et al., 2023a). Crowdsourcing work applies dynamic architectures to a wider variety of subjective creative tasks (Lasecki et al., 2015; Retelny et al., 2014; Chilton et al., 2019; Kittur et al., 2011) and includes more complex dynamic workflows such as multi-level map-reduce (Kittur et al., 2011) and dynamically choosing between workflows using their maximum expected value of utility (Goto et al., 2016). Decision theoretic techniques, such as Partially Observable Markov Decision Processes, yield great efficiencies in crowdsourcing (Dai et al., 2010, 2013), but have thus far barely been considered for LLM chaining (Madaan et al., 2023).
We recommend pursuing a broader range of subjective creative tasks with more communicative (e.g. establishing a hierarchy of roles) and dynamic (e.g. multi-level map-reduce) architectures.
@@ 9.2. Integrating the user
Although many chaining papers discuss involving humans into the chain hypothetically (Schick et al., 2022; Yang et al., 2022; Wu et al., 2023a; Wang et al., 2022; Wu et al., 2024), few implement this user oversight (Stretcu et al., 2023; Toubal et al., 2024). Our case studies showcase one example of interactions through directly manipulable outputs, but the varied ways of incorporating the user in crowdsourcing workflows could continue to inspire similar approaches in chaining.
Explorable interfaces surface disagreements to the user rather than resolving them in the workflow itself (Willett et al., 2012; Chang et al., 2017; Cheng and Bernstein, 2015; Mohanty et al., 2019). Worker-initiated guidance allows workers to define when to alert the user, focusing user attention only on what matters to the workers (Salehi et al., 2017; Kulkarni et al., 2012). In subjective tasks, quick iteration and choosing among uncertain options reduces the effort it takes for the user to prototype creative outputs and allows the user to guide ideation (Nebeling et al., 2016; Lasecki et al., 2015). For example, unstructured text can be interpreted in many ways depending on an analyst’s research questions, so recent work such as Concept Induction allows analysts to steer the process of LLM-aided text analysis (Lam et al., 2024).
We can also consider how our case study examples could be expanded to incorporate the user differently. With inspiration from the crowdsourcing workflow Revolt (Chang et al., 2017), a user of Cascade-LM could review edge cases that only match one category label of two that were merged. Soylent-LLM could replace the verify step with a user, as the user may recognize subtle undesirable changes in fixed phrases that are difficult to pre-define through prompts. Mechanical Novel-LLM could implement the quality thresholds strategy by having the user review after each iteration if the generation process should continue. The user could also give more oversight on the edit suggestions chosen. Incorporating the user in more ways reduces expense by calling LLMs fewer times and by directing the calls of LLMs that do occur towards more important needs. However, it does require more effort on the user’s part and could increase latency if the user is slower than the equivalent call to an LLM.
@@ 9.3. Studying chains more systematically
More systematic studies on optimal chain design would be valuable. Several studies already approach questions around model choice (Madaan et al., 2023; Du et al., 2023), incorporating roles into prompts (Wu et al., 2023a; Park et al., 2022), and using self-reflection or debate to validate outputs (Du et al., 2023; Liang et al., 2023). Although the outcomes may be different, systematic crowdsourcing studies could provide inspiration for some of these explorations. Prior crowdsourcing studies investigate the effect of context timing, architecture, and content on task output (Luther et al., 2015; Salehi et al., 2017). Others measure tradeoffs of iterative and parallel workflows (e.g., effect on latency (Luther et al., 2015), quality averages and variance (Little et al., 2010a; Kittur et al., 2011; Goto et al., 2016), risk of cascading errors (Goto et al., 2016)) in various task contexts, as well as good practices for each (e.g., appropriate number of branches or iterations (André et al., 2014; Huffaker et al., 2020)).
By specifying a strategy of interest, the design space can orient studies for understanding optimal tactical design decisions. For example, studying debate designs could look at debate to establish adaptability for tasks with many edge cases, debate to validate outputs, or debate to encourage a diversity of creative ideas for a subjective task. Evaluating the use of tactic designs within the context of specific strategies will lead to more generalizable findings than evaluating solely on task performance.